C3H8  + 5O2  --> 3CO2  + 4 H2O

Using the above reaction, calculate how many grams of H2O will be made if 195.0 g of O2 is reacted?


1
Expert's answer
2021-05-31T02:02:06-0400

MM(O2) = 32 g/mol

n(O2) "= \frac{195}{32}=6.093 \;mol"

According to the reaction:

n(H2O) "= \frac{4}{5}n(O_2) = \frac{4}{5} \times 6.093 = 4.875 \;mol"

MM(H2O) = 18 g/mol

m(H2O) "= 4.875 \times 18 = 87.75 \;g"

Answer: 87.75 g